(Please show calculations for study purposes)

Required information

[The following information applies to the questions displayed below.]

Cane Company manufactures two products called Alpha and Beta that sell for $225 and $175, respectively. Each product uses only one type of raw material that costs $6 per pound. The company has the capacity to annually produce 130,000 units of each product. Its average cost per unit for each product at this level of activity are given below:

Alpha Beta
Direct materials $ 42 $ 24
Direct labor 42 32
Variable manufacturing overhead 26 24
Traceable fixed manufacturing overhead 34 37
Variable selling expenses 31 27
Common fixed expenses 34 29
Total cost per unit $ 209 $ 173

The company considers its traceable fixed manufacturing overhead to be avoidable, whereas its common fixed expenses are unavoidable and have been allocated to products based on sales dollars.

Required:  Constrained Resources - Most Profitable Use Decision

a) How many pounds of raw material are needed to make one unit of each of the two products?

b) What contribution margin per pound of raw material is earned by each of the two products? (Round your answers to 2 decimal places.)

c) Assume that Cane’s customers would buy a maximum of 99,000 units of Alpha and 79,000 units of Beta. Also assume that the company’s raw material available for production is limited to 344,000 pounds. How many units of each product should Cane produce to maximize its profits?

d) Assume that Cane’s customers would buy a maximum of 99,000 units of Alpha and 79,000 units of Beta. Also assume that the company’s raw material available for production is limited to 344,000 pounds. What is the maximum contribution margin Cane Company can earn given the limited quantity of raw materials?
Answer #1

a) Pounds of raw materials are needed to make one unit

Alpha Beta
Direct material 42 24
Cost per pound 6 6
Pound per unit 7 4

b) Contribution margin per pound

Alpha Beta
Selling price 225 175
Variable cost 141 107
Contribution margin 84 68
Pound per unit 7 4
Contribution margin per pound 12 17

c) Optimum mix

Unit Pound
Beta 79000 79000*4 = 316000
Alpha 28000/7 = 4000 28000
Total 344000

d) Contribution margin = (79000*68+4000*84) = $5708000